1 First section

There should be hyperlinks for pages to follow the cursor.

Contents

1 First section  1
1.1 Sub . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1
1.1.1 SubSub . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2
2 Foo  4
2.1 Bar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

Some text in the section.

1.1 Sub

Some text in the subsection. Another subsection on next page.

1.1.1 SubSub Some text this next page in the sub-subsection.

Move cursor to see if pages follow cursor. aaa hjkhjk h

3 khjkhkjk hjkh kjh jk hjh kj hjkh k uu iuio uoi foo hjkhjkhk dfsf dddd hjh jkhjkhk The cursor should follow own other pages as well. This requires loading the

package hyperref, which is done automatically in pdf mode (if after loading the package ifpdf, the command \ifpdf is true).

2 Foo

2.1 Bar

Par And a paragraph.